# Attention as an Adaptive Filter

### v2

@@ B.4.1 Open loop waypoint control
B.4.2 Closed-form expressions for control and state evolution
@@ B.5 Extension to 2D Data
We again assume that all system and noise matrices are simultaneously diagonalizable: $\mathbf{A}_{\ell}=\mathbf{S}\bm{\Lambda}_{\ell}\mathbf{S}^{-1}$, $\mathbf{Q}_{\ell}=\mathbf{S}\bm{\Lambda}_{Q,\ell}\mathbf{S}^{-1}$, $\mathbf{R}=\mathbf{S}\bm{\Lambda}_{R}\mathbf{S}^{-1}$, for some invertible $\mathbf{S}\in\mathbb{C}^{d\times d}$, and diagonal $\bm{\Lambda}_{\ell},\bm{\Lambda}_{Q,\ell},\bm{\Lambda}_{R}$.
@@
$\sigma_{V}^{2}=\Big(\sigma_{x}^{2}\frac{e^{-2\alpha_{x}^{2}|\Delta x^{\prime}|}-1}{-2\alpha_{x}^{2}}\Big)\Big(\sigma_{y}^{2}\frac{e^{-2\alpha_{y}^{2}|\Delta y^{\prime}|}-1}{-2\alpha_{y}^{2}}\Big)+\eta^{2}e^{-2\alpha_{x}^{2}|\Delta x^{\prime}|-2\alpha_{y}^{2}|\Delta y^{\prime}|}$
@@
B.6 Further Justification of the Radial-Tangential Structure
B.6.1 Normal plus Rank-1 Noise Covariance
@@
B.6.2 Expressiveness of Radial-Tangential SDEs

### v3

@@ B.5 Model Predictive Control
The reference waypoint $\mathbf{z}_{\mathbf{s},\mathrm{ref}}$ is generally unknown. We can represent it using e.g. a low-rank parameterization. Define $\mathbf{W}_{1}\in\mathbb{C}^{d\times d_{\text{low}}},\,\mathbf{W}_{2}\in\mathbb{C}^{d_{\text{low}}\times d}$, where $d_{\text{low}}\ll d$, and let:
$\mathbf{z}_{\mathbf{s},\mathrm{ref}}=\mathbf{W}_{1}\mathbf{W}_{2}(e^{\bm{\lambda}\Delta t}\odot\mathbf{\bar{z}}_{\mathbf{s}})$
@@ B.6 Extension to 2D Data
We again assume that all system and noise matrices are simultaneously-diagonalizable: $\mathbf{A}_{\ell}=\mathbf{S}\bm{\Lambda}_{\ell}\mathbf{S}^{-1}$, $\mathbf{Q}_{\ell}=\mathbf{S}\bm{\Lambda}_{Q,\ell}\mathbf{S}^{-1}$, $\mathbf{R}=\mathbf{S}\bm{\Lambda}_{R}\mathbf{S}^{-1}$, for some invertible $\mathbf{S}\in\mathbb{C}^{d\times d}$, and diagonal $\bm{\Lambda}_{\ell},\bm{\Lambda}_{Q,\ell},\bm{\Lambda}_{R}$.
@@
$\sigma_{V}^{2}=\Big(\sigma_{x}^{2}\frac{e^{-2\alpha_{x}^{2}|\Delta x^{\prime}|}-1}{-2\alpha_{x}^{2}}\Big)\Big(\sigma_{y}^{2}\frac{e^{-2\alpha_{y}^{2}|\Delta y^{\prime}|}-1}{-2\alpha_{y}^{2}}\Big)+\eta^{2}e^{-2\alpha_{x}^{2}|\Delta x^{\prime}|-2\alpha_{y}^{2}|\Delta y^{\prime}|}+\gamma^{2}$
@@
B.7 Further Justification of the Radial-Tangential Structure
B.7.1 Normal plus Rank-1 Noise Covariance
@@
B.7.2 Expressiveness of Radial-Tangential SDEs